OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Enter your answer as an ordered pair. y = 2x - 3 y = -x + 3

OpenStudy (anonymous):

2x-3=-x+3 3x=6 x=2 sub to y=-x+3 y=-2+3 =1 (2,1)